# Model Context Protocol (MCP) Tool Descriptions Are Smelly! Towards Improving AI Agent Efficiency with Augmented MCP Tool Descriptions

### v1

@@ Abstract.
The Model Context Protocol (MCP) standardizes how Foundation Model (FM)-based agents interact with external systems by invoking tools. However, to understand a tool’s purpose and features, FMs rely on natural-language tool descriptions, making these descriptions a critical component in guiding FMs to select the optimal tool for a given (sub)task and to pass the right arguments to the tool. While defects or smells in these descriptions can misguide FM-based agents, their prevalence and consequences in the MCP ecosystem remain unclear.
To address this, we conduct the first large-scale empirical study of 856 tools spread across 103 MCP servers, assessing their description quality and their impact on agent performance. We identify six components of tool descriptions from the literature, develop a scoring rubric utilizing these components, then formalize tool description smells based on this rubric. By operationalizing this rubric through an FM-based scanner, we find that 97.1% of the analyzed tool descriptions contain at least one smell, with 56% failing to state their purpose clearly. While augmenting these descriptions for all components improves task success rates by a median of 5.85 percentage points and improves partial goal completion by 15.12%, it also increases the number of execution steps by 67.46% and regresses performance in 16.67% of cases. These findings highlight a trade-off between agent performance and cost, as well as the context sensitivity of the performance gain. Furthermore, component ablations show that compact variants of different component combinations often preserve behavioral reliability while reducing unnecessary token overhead, enabling more efficient use of the FM context window and lower execution costs.
@@
copyright: acmcopyrightjournal: TOSEMpublicationmonth: 0journalyear: TBDccs: Software and its engineering Empirical software validation
@@ 1. Introduction
In MCP-enabled workflows with these official and community MCP servers, tool descriptions serve as the primary semantic interface that guides FM behavior. These descriptions convey a tool’s intended functionality, constraints, and usage cues, and shape tool selection, parameterization, and multi-step orchestration (Hou et al., 2025). For example, as illustrated in Figure 1, upon receiving a user query (e.g., “What was the third-quarter income of Apple for 2025?”), the agent retrieves the names, descriptions, and schemas of all available tools from the connected MCP servers and injects this metadata into the FM’s context, along with the user query. Only then can the FM leverage these artifacts to discover the capabilities of the available tools, plan a solution strategy, select an appropriate tool (get_financial_statement), infer the required parameters (ticker="AAPL", financial_type="quarterly_income_statement"), and issue a tool call via the agent’s MCP client. The agent then returns the tool response to the FM, which synthesizes the final answer for the user.
@@
On the other hand, the efficiency of MCP-enabled agents has recently come under strict scrutiny, as the tool metadata, repeatedly injected into the FM’s context during a typical interaction with an FM-based agent, is inflating token usage and increasing execution cost. While emerging techniques like Agent Skills (PBC, 2025a) attempt to complement the MCP’s use by progressive discovery of capabilities to an FM, and Tool Search (PBC, 2025e) is enabling the FM to search for a tool on demand rather than always loading them into context, they do not eliminate the reliance on MCP tools and their tool descriptions. Instead, these developments highlight a fundamental tension in MCP-enabled agents: resolving tool description smells by augmenting all components may improve semantic guidance and agent performance, but it also consumes scarce context windows and increases costs. Any attempt to augment tool descriptions must therefore justify its cost and, ideally, identify compact representations that preserve effectiveness.
@@
To bridge this gap, we conduct the first large-scale empirical study of MCP tool description quality and its effect on FM-based agent performance. On a dataset of 103 major MCP servers comprising 856 tools, we scan tool descriptions using a structured quality rubric designed to identify potential suboptimal design patterns or smells (Moha et al., 2009). We then use FMs to automatically fix the identified smells and augment the description. Finally, we assess the impact of these augmented descriptions on FM-based agent performance using the MCP Universe benchmark (Luo et al., 2025). Our analysis aims to answer the following Research Questions (RQs).
@@
The primary contributions of this study are as follows:
@@ 3.1. Model Context Protocol (MCP)
To let AI applications operate on real systems, FM-based agents rely on tools that perform external actions, e.g., web search, database queries, API calls, code execution, or device control. Anthropic proposed the Model Context Protocol (MCP) to standardize how agents discover and invoke such tools via a common client-server protocol, reducing custom glue code across models and frameworks (Anthropic, 2025). MCP is positioned as an open standard for connecting AI applications to external systems, analogous to a universal port for agents. Since its introduction, the protocol has been adopted by major FM providers such as OpenAI, Microsoft, Google, and Cloudflare (Hasan et al., 2025b), and currently observes over 20 million weekly downloads for the Python and JavaScript SDKs of MCP servers, signaling strong community acceptance and usage.
@@
These descriptions play a central role at runtime, shaping how agents reason about and execute tool calls. As shown in Figure 1, MCP standardizes how FM-based agents interact with tools through a client-server loop. The MCP client mediates between the FM and one or more MCP servers, each exposing tool metadata, e.g., name, description, and input schema, to the model.
@@
Given this neatly coupled reasoning loop between the FM and tool descriptions, evaluating MCP-enabled agents requires benchmarks that can faithfully capture both planning correctness and execution behavior. In this study, we adopt the MCP-Universe benchmark (Luo et al., 2025), which is one of the most comprehensive and widely used benchmarks for evaluating MCP-based agents. MCP-Universe spans multiple domains such as finance, data analysis, repository management, and information retrieval, and defines realistic goal-oriented tasks that test an agent’s ability to reason, plan, and execute MCP tools effectively. It provides a total of 18 MCP servers with 202 tools combined. Each task is evaluated by at least one evaluator, with an average of 3.3 evaluators per task.
@@ 3.2. Studies on Smells
Within the MCP ecosystem, recent work reports code smells in MCP servers (Hasan et al., 2025b). To the best of our knowledge, no empirical study has examined design-level or specification-level smells for MCP servers. Our study addresses this gap by focusing on tool descriptions as a first-class artifact and by analyzing the smells hidden in these tool descriptions.
@@ 3.3. Refactoring the Smells and Optimization
Beyond static analysis, a growing body of work has explored machine learning (ML) and deep learning (DL) techniques to identify refactoring opportunities and predict refactoring actions. These include multi-layer perceptrons (MLPs) and recurrent neural network (RNN) variants such as bidirectional long short-term memory (BiLSTM) and gated recurrent unit networks (GRU) for refactoring type prediction (Mohan et al., 2016; Szalontai et al., 2021), convolutional neural network (CNN)- and RNN-based models with embedding pipelines for naming-related refactorings such as Rename Method (Liang et al., 2021), and RNN encoder-decoder architectures that model refactoring as a learned code transformation task (Tufano et al., 2019), as summarized in the survey by Naik et al. (Naik et al., 2024).
@@
Despite the functional similarities between prompts and tool descriptions in the MCP ecosystem, no prior study has investigated the optimization or refactoring of tool descriptions. Our work addresses this gap by examining how FM-based optimization can be adapted to improve the quality of MCP tool descriptions.
@@ 4. Methodology
In this section, we present our study design as outlined in Figure 3.
Figure 3. Overview of the study design. The components in bright yellow boxes represent workflows repurposed from the MCP-Universe benchmark for evaluation and benchmarking, while all other components and processes were introduced in this study.
@@ 4.2.2. Tool description extraction via MCP client
To extract the tool descriptions from MCP servers, we develop a lightweight MCP client. The MCP client communicates with MCP servers dynamically through the reflection protocol (Hasan et al., 2025b) by sending a tools/list request. In response, each server returns a structured array containing all available tools along with their names, descriptions, and input schema. We implement this dynamic approach, avoiding static analysis, to ensure generalizability across the heterogeneous MCP ecosystem. Unlike static analysis, which relies on source code availability of MCP servers and language-specific parsers, this approach utilizes the native reflection of MCP to extract descriptions from any MCP server, regardless of its underlying implementation language (e.g., Python, Rust) or proprietary ownership status (e.g., open source or proprietary). The scanner then extracts the tool descriptions and input schema from the MCP client and prepares them for use in the subsequent evaluation step.
@@ 4.3.3. Smell identification
To transition from the raw multi-model scores to definitive smell assignments, we apply an aggregation and thresholding procedure. For each tool description and each component, we compute the arithmetic mean of the three scores assigned independently by the FM evaluators. Formally, for a component with scores Score1, Score2, Score3:
@@ 4.4. Resolving Smells via Tool Component Augmentation
The objective of tool description augmentation is to fix the smells identified in Section 4.3 while preserving the semantic fidelity of the original content. To achieve this, we design a semi-automated augmentor that combines rubric-based augmentation with an FM to generate refined, comprehensive, and factually consistent tool descriptions, being motivated by prior FM-based prompt and text optimization techniques (Opsahl-Ong et al., 2024; Khattab et al., 2023).
@@ 4.4.1. Initial augmentation of components
In the first stage, the augmentor collects the original tool descriptions of all the MCP servers used in the MCP-Universe benchmark (Luo et al., 2025) and the input schema from MCP servers through an MCP client following the same procedure described in Section 4.2.2. It then applies the rubrics defined in Section 4.1 to guide the augmentation process. Using the GPT-4.1-mini model as FM, the augmentor automatically enhances each description by improving coverage across five rubric components: Purpose, Guidelines, Limitations, Parameter Explanation, and Length. The outputs from this stage are referred to as init_augmented_description, as these outputs serve as input for the next refinement step. We intentionally exclude the Examples component at this stage because the model cannot reliably generate factually grounded examples without execution traces, and doing so would risk introducing hallucinated or incorrect examples. This issue is addressed in the next stage, where examples are constructed from actual tool executions.
@@ 5.1. RQ-1: To what extent do MCP tools’ descriptions contain smells?
Approach. We develop a rubric for evaluating MCP tool descriptions which consists of the components of tool description derived from Anthropic’s design guidelines, practitioners’ recommendations, and prior research, as well as a structured scoring scale as mentioned in Section 4.1. As FMs are the primary consumers of the tool description, to evaluate whether FMs can understand and interpret these tool descriptions, we score the 856 tools collected from MCP servers through an FM-based scanner. This scanner consists of an LLM-as-Jury with 3 FMs using the rubric, as detailed in Section 4.3. Then we identify the cases with low scores and map those to recurring smells. Given prior work showing differences between official and community-maintained MCP servers (Hasan et al., 2025b), we further compare these scores statistically using pairwise Mann-Whitney U tests with Bonferroni correction (Ruxton and Beauchamp, 2008).
@@
Figure 4 (presented in Section 4.1.3) provides an example of a high-quality tool description from the Sequential Thinking MCP server. The description clearly conveys the tool’s purpose, guidelines, limitations, and parameters. In contrast, Figure 9 shows three examples of low-quality tool descriptions that scored below the median quality score. These minimal descriptions provide little contextual information, leaving foundation models with insufficient cues to infer when or how to use the tools effectively.
@@
In addition, we test whether the observed changes in SR, AE, and AS are statistically significant. As the original study did not provide these metrics for every (task, model) combination, we measure these metrics using the model introduced in our work, i.e., Qwen3-Next-80B-A3B-Instruct. Since SR is a binary outcome (0/1), we assess the significance of SR changes using McNemar’s test in its chi-squared formulation (Pembury Smith and Ruxton, 2020). In contrast, as AE and AS are continuous-valued metrics, we evaluate their statistical significance using the Wilcoxon signed-rank test.
@@
Beyond task-level success rate, augmented tool descriptions also improve evaluator-level performance, increasing the Average Evaluator Score (AE) by 15.12% across all foundation models. As shown in Table 6, the overall AE, which represents the mean proportion of evaluators passing per task (where AE = 1 denotes complete success) increased consistently across all four models. This indicates that even for tasks not fully completed (i.e., not satisfying all evaluators), augmented descriptions help agents satisfy a greater share of evaluation criteria. The Wilcoxon signed-rank test confirms that this improvement in AE is statistically significant (with $p<0.01$). Furthermore, we observe that 17 tasks, or 7.36% of all tasks (median across models), can achieve an AE $\geq 0.80$ (i.e., passing more than 80% of evaluators) yet still fall short of final success, often because the benchmark imposes a maximum iteration limit. These trends suggest that augmentation enhances intermediate reasoning and partial goal completion, aligning with prior findings (Sahoo et al., 2024; Zheng et al., 2023) that improved prompt clarity strengthens reasoning pathways to drive steadier, more progressive task execution.
@@
Findings. There is no single “golden” combination of components that yields the best results across all domain-model pairs. As summarized in Table 7, the best-performing component combination varies across domain-model pairs. For instance, in the Finance domain with GPT-4.1, using only the Purpose and Guidelines components yields the highest success rate, surpassing the fully augmented tool description, making it the best-performing component combination (BC). In contrast, tasks in Location Navigation with the same model show the best performance with the fully augmented tool description (FR). Meanwhile, in Repository Management with Qwen3-Coder-480B-A35B, performance improves modestly when only the Examples component is removed. These findings indicate that the contribution of individual rubric components is context-dependent, influenced by both model architecture and domain characteristics.

### v2

@@ Abstract.
The Model Context Protocol (MCP) introduces a standard specification that defines how Foundation Model (FM)-based agents should interact with external systems by invoking tools. However, to understand a tool’s purpose and features, FMs rely on natural-language tool descriptions, making these descriptions a critical component in guiding FMs to select the optimal tool for a given (sub)task and to pass the right arguments to the tool. While defects or smells in these descriptions can misguide FM-based agents, their prevalence and consequences in the MCP ecosystem remain unclear.
Hence, we examine 856 tools spread across 103 MCP servers empirically, assess their description quality, and their impact on agent performance. We identify six components of tool descriptions from the literature, develop a scoring rubric utilizing these components, and then formalize tool description smells based on this rubric. By operationalizing this rubric through an FM-based scanner, we find that 97.1% of the analyzed tool descriptions contain at least one smell, with 56% failing to state their purpose clearly. While augmenting these descriptions for all components improves task success rates by a median of 5.85 percentage points and improves partial goal completion by 15.12%, it also increases the number of execution steps by 67.46% and regresses performance in 16.67% of cases. These results indicate that achieving performance gains is not straightforward; while execution cost can act as a trade-off, execution context can also impact. Furthermore, component ablations show that compact variants of different component combinations often preserve behavioral reliability while reducing unnecessary token overhead, enabling more efficient use of the FM context window and lower execution costs.
@@
copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Software and its engineering Empirical software validation
@@ 1. Introduction
In MCP-enabled workflows with these official and community MCP servers, tool descriptions serve as the essential linguistic guide that directs FM behavior. These descriptions convey a tool’s intended functionality, constraints, and usage cues, and shape tool selection, parameterization, and multi-step orchestration (Hou et al., 2025). For example, as illustrated in Figure 1, upon receiving a user query (e.g., “What was the third-quarter income of Apple for 2025?”), the agent retrieves the names, descriptions, and schemas of all available tools from the connected MCP servers and injects this metadata into the FM’s context, along with the user query. Only then can the FM leverage these artifacts to discover the capabilities of the available tools, plan a solution strategy, select an appropriate tool (get_financial_statement), infer the required parameters (ticker="AAPL", financial_type="quarterly_income_statement"), and issue a tool call via the agent’s MCP client. The agent then returns the tool response to the FM, which synthesizes the final answer for the user.
@@
On the other hand, the efficiency of AI agents using MCP servers has recently come under strict scrutiny, as the tool metadata, repeatedly injected into the FM’s context during a typical interaction with an FM-based agent, is inflating token usage and increasing execution cost. While emerging techniques like Agent Skills (PBC, 2025a) attempt to complement the MCP’s use by progressive discovery of capabilities to an FM, and Tool Search (PBC, 2025e) is enabling the FM to search for a tool on demand rather than always loading them into context, they do not eliminate the reliance on MCP tools and their tool descriptions. Instead, these developments highlight a fundamental tension in MCP-enabled agents: resolving tool description smells by augmenting all components may improve semantic guidance and agent performance, but it also consumes scarce context windows and increases costs. Any attempt to augment tool descriptions must therefore justify its cost and, ideally, identify compact representations that preserve effectiveness.
@@
Recognizing this lack of systematic research,, we carry out an extensive empirical study of MCP tool description quality and its effect on FM-based agent performance. On a dataset of 103 major MCP servers comprising 856 tools, we scan tool descriptions using a structured quality rubric designed to identify potential suboptimal design patterns or smells (Moha et al., 2009). We then use FMs to automatically fix the identified smells and augment the description. Finally, we utilize the MCP Universe benchmark (Luo et al., 2025) to evaluate how these augmented descriptions influence FM-based agents’ performance. This evaluation is structured around the following core Research Questions (RQs).
@@
Our study makes the following key contributions to MCP ecosystem:
@@ 3.1. Model Context Protocol (MCP)
To let AI applications operate on real systems, FM-based agents rely on tools that perform external actions, e.g., searching the web for information retrieval, executing database queries, triggering API calls, running programs, or performing device management. Anthropic proposed the Model Context Protocol (MCP) so that AI agents can discover tools using a standard client-server protocol and invoke them with the proper arguments, reducing the need for custom glue code across models and frameworks. Since then, MCP has been evolving as an open integration protocol that facilitates communication between AI applications and external systems, often positioned as a “USB-C port for AI” (Anthropic, 2025). Since its introduction, the protocol has been adopted by major FM providers such as OpenAI, Microsoft, Google, and Cloudflare (Hasan et al., 2025b), and currently observes over 20 million weekly downloads for the Python and JavaScript SDKs of MCP servers, signaling strong community acceptance and usage.
@@
These descriptions play a central role at runtime, shaping how agents reason about and execute tool calls. As shown in Figure 1, MCP standardizes how FM-based agents interact with tools through a client-server loop. The MCP client orchestrates this communication loop among the FM and available MCP servers, each of which exposes tool metadata, e.g., name, description, and input schema, to the model.
@@
Given this neatly coupled reasoning loop between the FM and tool descriptions, evaluating MCP-enabled agents requires benchmarks that can faithfully capture both planning correctness and execution behavior. In this study, we adopt the MCP-Universe benchmark (Luo et al., 2025), which is one of the most comprehensive and widely used benchmarks for evaluating MCP-based agents. MCP-Universe spans multiple domains, such as finance, data analysis, repository management, and information retrieval, and defines realistic, goal-oriented tasks that test an agent’s ability to choose the correct tool and execute it with the proper arguments. It provides a total of 18 MCP servers with 202 tools combined. Each task is evaluated by at least one evaluator, with an average of 3.3 evaluators per task.
@@ 3.2. Studies on Smells
Within the MCP ecosystem, recent work reports code smells in MCP servers (Hasan et al., 2025b). However, to date, no empirical analyses have examined their design- or specification-level “smells“. In this study, we focus on tool descriptions as a first-class artifacts and, by analyzing the smells hidden within them, we provide empirical evidence on the impact of tool descriptions on agent performance.
@@ 3.3. Refactoring the Smells and Optimization
Beyond static analysis, a growing body of work has operationalized different machine learning (ML) and deep learning (DL) based techniques to identify refactoring opportunities and predict refactoring actions. These include multi-layer perceptrons (MLPs) and recurrent neural network (RNN) variants such as bidirectional long short-term memory (BiLSTM) and gated recurrent unit networks (GRU) for refactoring type prediction (Mohan et al., 2016; Szalontai et al., 2021), convolutional neural network (CNN)- and RNN-based models with embedding pipelines for naming-related refactorings such as Rename Method (Liang et al., 2021), and RNN encoder-decoder architectures that model refactoring as a learned code transformation task (Tufano et al., 2019), as summarized in the survey by Naik et al. (Naik et al., 2024).
@@
Despite the functional similarities between prompts and tool descriptions in the MCP ecosystem, no prior study has investigated the optimization or refactoring of tool descriptions. To fill this gap, we examine how FM-based augmentation can be adapted to improve the quality of MCP tool descriptions.
@@ 4. Methodology
We illustrate the methodology used in this study in Figure 3 and describe each step in the following sub-sections:
Figure 3. Overview of the process used to study MCP server tool descriptions. The components in bright yellow boxes represent workflows repurposed from the MCP-Universe benchmark for evaluation and benchmarking, while all other components and processes were introduced in this study.
@@ 4.2.2. Tool description extraction via MCP client
We develop a lightweight MCP client to extract the tool descriptions from MCP servers. This MCP client dynamically communicates with MCP servers through the reflection protocol (Hasan et al., 2025b) by sending a tools/list request. In response, each server returns a structured array containing all available tools along with their names, descriptions, and input schema. We implement this dynamic approach, avoiding static analysis, to ensure generalizability across the heterogeneous MCP ecosystem. Unlike static analysis, which relies on source code availability of MCP servers and language-specific parsers, this approach utilizes the native reflection of MCP to extract descriptions from any MCP server, regardless of its underlying implementation language (e.g., Python, Rust) or proprietary ownership status (e.g., open source or proprietary). The scanner then extracts the tool descriptions and input schema from the MCP client and prepares them for use in the subsequent evaluation step.
@@ 4.3.3. Smell identification
To transition from the raw multi-model scores to definitive smell assignments, we apply an aggregation and thresholding procedure. For each tool description and each component, we calculate the average of the three scores given by the FM evaluators and consider this as the consolidated score. Formally, for a component with scores Score1, Score2, Score3:
@@ 4.4. Resolving Smells via Tool Component Augmentation
The objective of the tool description augmentation is to fix the smells identified in Section 4.3 while retaining the original meaning and intent of the tool description. To achieve this, we design a semi-automated augmentor that combines rubric-based augmentation with an FM to generate refined, comprehensive, and factually consistent tool descriptions, being motivated by prior FM-based prompt and text optimization techniques (Opsahl-Ong et al., 2024; Khattab et al., 2023).
@@ 4.4.1. Initial augmentation of components
In the first stage, the augmentor collects the original tool descriptions from every MCP server included in the MCP-Universe (Luo et al., 2025) and the input schema from MCP servers through an MCP client following the same procedure described in Section 4.2.2. It then applies the rubrics defined in Section 4.1 to guide the augmentation process. Using the GPT-4.1-mini model as FM, the augmentor automatically enhances each description by improving coverage across five rubric components: Purpose, Guidelines, Limitations, Parameter Explanation, and Length. The outputs from this stage are referred to as init_augmented_description, as these outputs serve as input for the next refinement step. We intentionally exclude the Examples component at this stage because the model cannot reliably generate factually grounded examples without execution traces, and doing so would risk introducing hallucinated or incorrect examples. This issue is addressed in the next stage, where examples are constructed from actual tool executions.
@@ 5.1. RQ-1: To what extent do MCP tools’ descriptions contain smells?
Approach. We develop a rubric for evaluating MCP tool descriptions which consists of the components of tool description derived from Anthropic’s design guidelines, practitioners’ recommendations, and prior research, as well as a structured scoring scale as mentioned in Section 4.1. As FMs are the primary consumers of the tool description, to evaluate whether FMs can understand and interpret these tool descriptions, we score the 856 tools collected from MCP servers through an FM-based scanner. This scanner consists of an LLM-as-Jury with 3 FMs using the rubric, as detailed in Section 4.3. Then we identify the cases with low scores and map those to recurring smells. Given prior work showing differences between official and community-maintained MCP servers (Hasan et al., 2025b), we conduct pairwise nonparametric comparisons using the Mann–Whitney U test and refine the results with Bonferroni-adjusted $p$-values (Ruxton and Beauchamp, 2008) to uncover any such differences.
@@
We illustrate a high-quality description of a tool in Figure 4 (presented in Section 4.1.3) from the Sequential Thinking MCP server. The description clearly conveys the tool’s purpose, guidelines, limitations, and parameters. In contrast, Figure 9 shows three examples of low-quality tool descriptions that scored below the median quality score. These minimal descriptions provide little contextual information, leaving foundation models with insufficient cues to infer when or how to use the tools effectively.
@@
In addition, we test whether the observed changes in SR, AE, and AS are statistically significant. As the original study did not provide these metrics for every (task, model) combination, we measure these metrics using the model introduced in our work, i.e., Qwen3-Next-80B-A3B-Instruct. Since SR is a binary outcome (0/1), we assess the significance of SR changes using McNemar’s test in its chi-squared formulation (Pembury Smith and Ruxton, 2020). In contrast, as AE and AS are continuous-valued metrics, we apply a Wilcoxon signed-rank test to determine their statistical significance.
@@
Beyond task-level success rate, augmented tool descriptions also improve evaluator-level performance, increasing the Average Evaluator Score (AE) by 15.12% across all foundation models. As shown in Table 6, the overall AE, which represents the mean proportion of evaluators passing per task (where AE = 1 denotes complete success) increased consistently across all four models. This indicates that even for tasks not fully completed (i.e., not satisfying all evaluators), augmented descriptions help agents satisfy a greater share of evaluation criteria. Using the Wilcoxon signed-rank procedure, we determine that the observed increase in AE is statistically significant, with $p<0.01$. Furthermore, we observe that 17 tasks, or 7.36% of all tasks (median across models), can achieve an AE $\geq 0.80$ (i.e., passing more than 80% of evaluators) yet still fall short of final success, often because the benchmark imposes a maximum iteration limit. These trends suggest that augmentation enhances intermediate reasoning and partial goal completion, aligning with prior findings (Sahoo et al., 2024; Zheng et al., 2023) that improved prompt clarity strengthens reasoning pathways to drive steadier, more progressive task execution.
@@
Findings. There is no single “golden” combination of components that yields the best results across all domain-model pairs. As summarized in Table 7, the best-performing component combination varies across domain-model pairs. For instance, in the Finance domain with GPT-4.1, using only the Purpose and Guidelines components yields the highest success rate, surpassing the fully augmented tool description, making it the best-performing component combination (BC). In contrast, tasks in Location Navigation with the same model show the best performance with the fully augmented tool description (FR). Meanwhile, in Repository Management with Qwen3-Coder-480B-A35B, performance improves modestly when only the Examples component is removed. These results reveals a complex interplay where the effectiveness of specific rubric components varies across settings, depending on both the underlying model architecture and the domain-specific requirements.